the edge length of FCC is 508 pm if the radius if the cation is 110 pm the radius of anion is           ans is 114 pm

Grade:12

3 Answers

siddharth gupta
9 years ago
the same question was asked in aieee 2010.The language of question is not very clear as FCC can be rock salt,zinc blende,fluorite or anti-fluorite , but we simply consider it to be rock salt 
and hence rc+ra=a/2i.e110+ra=254i.e ra=144 pm.
thank you
siddharth gupta
Sai lokheshh
6 years ago
Rc+Ra=1/2,which implies that answer will 249 pico metre it can also be find by drawing the cube and filled by a pythagorous theorem which implies answer as 249 pico metre
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

FCC can be rock salt,zinc blende,fluorite or anti-fluorite , but we simply consider it to be rock salt
and hence rc+ra=a/2i.e110+ra=254i.e ra=144 pm.

Thanks and Regards